OpenStudy (anonymous):

Maria is now 4 times older than Kate. Four years ago, Maria was six times as old as Kate. Find each of their ages now.

OpenStudy (kaederfds):

M=4K M-4=6(K-4) 4K-4=6K-24 2K=20 K=10 M=4.10=40 Kate is 10 & Maria is 40

OpenStudy (anonymous):

Oh thank you :D I think that i understand it now